A line passes through (-4,5) with an x-intercept of 4. Write the equation in slope-intercept form.
Answer by Alan3354(69443)

A line passes through (-4,5) with an x-intercept of 4. Write the equation in slope-intercept form.
--------------
An x-intercept of 4 is the point (4,0).
Find the slope, m, of the line thru the 2 points.
m = diffy/diffx = (5-0)/(-4-4)
m = -5/8
---------
Then use
y-y1 = m*(x-x1) where (x1,y1) is either point.
y-0 = (-5/8)*(x-4)
y = (-5/8)x + 5/2